OpenStudy (clamin):

Please help!! Medal! Hiw can I solve this?? f(x)=8x^3+14x^2+11x+3

OpenStudy (clamin):

Fund the number of zeros of the polynomial function.

Vocaloid (vocaloid):

hint: what is the highest exponent?

OpenStudy (clamin):

3

OpenStudy (clamin):

@vocaloid

OpenStudy (anonymous):

derivatives.

OpenStudy (amistre64):

rational root thrm might prove useful as well

OpenStudy (amistre64):

real zeros, or zeros in general?

OpenStudy (clamin):

Rational zeros??

OpenStudy (amistre64):

rational numbers are real numbers ...

OpenStudy (clamin):

Teach me how to solve this cuz i never solve like this equation before

OpenStudy (amistre64):

the number of zeroes is determined by the degree of the polynomial the value of the zeroes is a different story

OpenStudy (mertsj):

Did you get it?

OpenStudy (anonymous):

\[f \left( x \right)=8x^3+14x^2+11x+3\] \[=8x^3+4x^2+10x^2+5x+6x+3\] \[=4x^2(2x+1)+5x(2x+1)+3(2x+1)\] \[=(2x+1)( 4x^2+5x+3) \]